OpenStudy (anonymous):

Expand ln((7(x+3)^5)/((x)^1/2))

OpenStudy (anonymous):

You can use all the laws for logarithm ln((7(x+3)^5)/((x)^1/2))=ln (7(x+3)^5) - ln (x)^1/2)= =ln7 +ln(x+3)^5 -1/2lnx=ln7+5ln(x+3)-1/2lnx